In Industry 1, Firm A, B, C, D and E each has 20 percent of the market. In Industry 2, Firm A has 75 percent of the market, Firm B has 2 percent, Firm C has 2 percent, and Firm D and 20 other firms all have 1 percent each. Which of these two industries do you consider likely to be more competitive, and why?

*step by step please*
Answer #1

To measure the competitiveness, we need to calculate the HHI for every industry.

In industry 1, the HHI for every firm will be the same. So the total HHI = 5* 20*20 = 2000

In industry 2, the HHI for firm A = 75*75 = 5625

HHI for firm B and C each = 2*2 = 4

HHI for firm D = 20*20 = 400

HHI for firm E = 1*1 = 1

Total HHI for industry B = 5625+4+4+400+1 = 6034.

So industry 1 is more competitive as its HHI is lesser than industry 2’s HHI.